Equal masses of paraffin and water are mixed in a container of negligible
thermal capacity. Initial temperature of water is 80
1
Expert's answer
2021-10-02T17:34:47-0400

"Q_1=Q_2,"

"c_1m_1(80-t)=\lambda m_2+c_2m_2(t-20),"

"m_1=m_2,"

"t=\frac{80c_1+20c_2-\lambda}{c_1+c_2}=53°C."